# Discrete Solution Operator Learning for Geometry-Dependent PDEs

### v1

@@ Abstract
Neural operator learning accelerates PDE solution by approximating operators as mappings between continuous function spaces. Yet in many engineering settings, varying geometry induces discrete structural changes, including topological changes, abrupt changes in boundary conditions or boundary types, and changes in the effective computational domain, which break the smooth-variation premise. Here we introduce Discrete Solution Operator Learning (DiSOL), a complementary paradigm that learns discrete solution procedures rather than continuous function-space operators. DiSOL factorizes the solver into learnable stages that mirror classical discretizations: local contribution encoding, multiscale assembly, and implicit solution reconstruction on an embedded grid, thereby preserving procedure-level consistency while adapting to geometry-dependent discrete structures. Across geometry-dependent Poisson, advection–diffusion, linear elasticity, as well as spatiotemporal heat-conduction problems, DiSOL produces stable and accurate predictions under both in-distribution and strongly out-of-distribution geometries, including discontinuous boundaries and topological changes. These results highlight the need for procedural operator representations in geometry-dominated regimes and position discrete solution operator learning as a distinct, complementary direction in scientific machine learning.
@@ 1 Introduction
The numerical solution of partial differential equations (PDEs) lies at the core of scientific computing and engineering analysis [2, 26, 1, 18]. For decades, mature discrete numerical algorithms, such as finite element [26], finite volume [14], and finite difference [29] methods, have provided reliable and accurate solutions across a wide range of applications. A defining characteristic of these methods is that they operate on discrete computational representations: local contributions are evaluated on elements, assembled into global representations, and subsequently solved through well-defined procedures. While geometry variation alters the instantiated discrete system, the underlying solution procedures by which local contributions are evaluated and assembled remain unchanged (Fig. 1a). That is, the same local evaluation and assembly rules are reused wherever they are activated, even though the activated locations change with geometry. From a functional perspective, these procedures define mappings from problem specifications to solutions, but are implemented through explicit local-to-global procedures.
Recently, data-driven surrogate and operator learning methods have emerged as a promising alternative for accelerating PDE solutions across varying inputs and physical parameters [5, 24, 3, 8]. Two complementary directions have been widely explored. One treats PDE learning as field-to-field regression on discrete grids (or strong discretization-based surrogates) using convolutional architectures, often in encoder-decoder or UNet-based forms [21, 20, 35], achieving strong empirical performance when inputs and outputs are naturally represented on fixed-resolution discretizations. The other focuses on continuous neural operators—such as DeepONet [22], the Fourier Neural Operator (FNO) [17], and their variants [9, 12, 13, 15, 34, 16, 10, 23, 31, 30, 7]—that approximate smooth mappings between input and output functions in continuous function spaces (Fig. 1d). Many of these methods have demonstrated impressive performance on benchmark problems with regular domains and smoothly varying geometries [32, 11]. These advances also highlight the role of inductive bias, i.e., structural assumptions encoded in model architectures or learning objectives that shape what solution mechanisms can be learned efficiently from finite data.
@@
This observation highlights a distinction that is often overlooked: although classical discrete numerical solvers can be regarded as operators, learning a continuous operator is not equivalent to learning a discrete numerical solution procedure. Classical solvers do not approximate a single global mapping; instead, they execute a sequence of fixed procedures—local contribution evaluation, global assembly, and solution—which is invariant under geometry variation, even though the assembled global information changes (Figs. 1a,b) [19, 27, 6, 4]. Approximating a smooth functional mapping may therefore be misaligned with the procedural nature of geometry-dependent solution processes.
@@
We validate DiSOL on four classes of geometry-dependent PDEs with increasing complexity: a Poisson equation, an advection–diffusion equation spanning diffusion- and transport-dominated cases, a vector-valued linear elasticity problem, and a spatiotemporal thermal conduction problem. Across these cases, DiSOL delivers stable and accurate predictions under both in-distribution (ID) and out-of-distribution (OOD) computational domains, whereas continuous neural operators exhibit structural degradation when geometry-induced discrete variation dominates the problem setting. Moreover, our results indicate that these failures arise primarily from a mismatch between the assumptions of continuous operators and geometry-induced discrete solution structures, rather than from optimization difficulty or insufficient model capacity.
@@
Fig. 1: Discrete Solution Operator Learning (DiSOL) in contrast to continuous neural operator paradigms. a, Classical discrete numerical methods for partial differential equations (PDEs), exemplified by the finite element method. A PDE operator can be realized through a sequence of fixed discrete procedures, including local operator evaluation (e.g., element-level stiffness and force computation), global assembly, and solution via assembled global systems. Under geometry variation, the assembled results change, whereas the underlying assembly procedures remain unchanged. b, An abstracted view of discrete problem solving, highlighting the canonical workflow shared by many numerical methods: local contributions evaluation, global information assembly, and solution through globally aggregated representations. c, Discrete Solution Operator Learning (DiSOL). Instead of explicitly constructing numerical matrices, DiSOL represents the discrete solution process through learnable neural operators that conceptually correspond to the local operator, the multiscale assembly operator, and the implicit problem-solving operator. These components collectively form a discrete operator that preserves procedure-level invariance under geometry variation, in the sense that the underlying solution procedure remains unchanged, while adapting its outputs to geometry-dependent discrete structures. d, Representative continuous neural operator paradigms, illustrated by DeepONet [22] and FNO [17]. These approaches model PDE operators as global mappings in continuous function spaces, using global basis combinations or spectral transformations. Such representations are primarily designed for smooth function-space mappings, and may become misaligned when geometry-induced domain changes and discontinuities dominate the problem setting. In contrast, DiSOL does not aim to approximate continuous operators directly, but learns discrete solution procedures, leading to a fundamentally different inductive bias in geometry-dependent problem settings.
@@ 2 Results
where $\Omega_{h}$ denotes the discrete computational domain (geometry mask and adjacency implied by the grid), $\Gamma_{h}$ encodes boundary-type and boundary value specifications on that domain, $f_{h}$ denotes source fields, $\eta_{h}$ collects problem-dependent physical parameters (e.g., diffusion coefficient, advection velocity, elastic moduli) that are fixed or prescribed for a given task, and $u_{h}$ denotes the normalized solution pattern field on the grid. In the main text we focus on learning the normalized solution pattern $u_{h}:=U_{h}/u_{\lim}$ (see Methods, Section 4.1, and Supplementary Information D); all models predict $\hat{u}_{h}$, and all reported metrics are computed on $u_{h}$ restricted to $\Omega_{h}$. All experiments below follow Eq. 1.
@@ 2.1 Geometry-dependent Poisson problem
Fig. 2: Discrete solution operator learning (DiSOL) for a geometry-dependent 2D Poisson problem. a, Problem formulation and data representation. The input to the model consists of three discrete fields defined on a fixed Cartesian grid: a geometry mask, a boundary-condition selection map, and a source term. The model outputs the corresponding discrete solution pattern. b, Training and validation loss histories for DiSOL, DeepONet and the FNO under comparable model capacity and training settings. DiSOL converges faster and achieves substantially lower validation loss (All models are trained under the same optimizer/schedule and have comparable parameter counts, $\approx 0.13$M each; details in Supplementary Information C.4 and 5.). c, ID test results for two representative cases. DiSOL accurately predicts the ground-truth solution patterns, while DeepONet shows pronounced distortions in the normalized pattern values (all fields are normalized patterns), despite the overall smoothness of the solutions. d, OOD generalization tests involving unseen geometries with sharp corners and internal holes, discontinuous and localized boundary conditions, and high-frequency source terms. DiSOL remains stable and consistent with the reference solutions, whereas DeepONet and FNO show substantial degradation in solution-pattern fidelity. See Supplementary Information A.1, B and E.1 for full setup and extended statistics.
@@
Figure 2c presents two representative ID test cases (more results are presented in Supplementary Information E.1). Although these samples match the training distribution in geometry and boundary configuration and yield globally smooth reference solutions, DeepONet already exhibits systematic deviations, most notably pattern mismatches and boundary-localized response errors near irregular geometric features. These deviations are structured rather than stochastic, indicating difficulty in capturing geometry-induced local contribution patterns within a continuous-operator parameterization. In contrast, DiSOL accurately reproduces both the global solution morphology and localized boundary responses.
@@
2.2 Geometry-dependent advection–diffusion problem
@@
Fig. 3: Discrete solution operator learning (DiSOL) results for the advection-diffusion equation. a,b, Representative ID test cases under two transport cases: a diffusion-dominated case ($\text{Pe}\approx 0.45$) and an advection-dominated case ($\text{Pe}\approx 4.5$). Ground-truth solutions are compared with predictions from the proposed discrete solution operator (DiSOL) and the FNO. As transport effects become stronger and solution structures align more closely with flow direction and complex geometric boundaries, DiSOL consistently preserves the global transport patterns and boundary-induced features, whereas FNO exhibits increasingly pronounced structural deviations, particularly near irregular geometric features. c,d, Training and validation loss histories for DiSOL, FNO, and DeepONet under the same two Péclet number settings ($\text{Pe}\approx 0.45$ and $\text{Pe}\approx 4.5$). In both cases, DiSOL converges rapidly and achieves substantially lower training and validation errors, while the continuous neural operator baselines saturate at significantly higher error levels. e, OOD generalization results for the advection-diffusion equation at $\text{Pe}\approx 4.5$, involving unseen geometries with sharp corners and topological changes, together with localized and discontinuous boundary conditions. DiSOL preserves the overall transport structure and solution morphology. f, Statistical comparison of relative L1 errors across different models under ID and OOD settings for both $\text{Pe}\approx 0.45$ and $\text{Pe}\approx 4.5$. Box plots (logarithmic scale) show that DiSOL consistently outperforms continuous neural operator baselines in both cases, and that performance degradation is primarily driven by geometric distribution shifts rather than by increased transport dominance. See Supplementary Information A.2, B and E.2 for full setup and extended statistics.
@@
We next consider a steady advection–diffusion problem to examine geometry generalization under transport-dominated dynamics. Although the governing equation and physical parameters are fixed within each case, variations in domain geometry induce sharp changes in effective transport pathways and boundary-layer structures. This case therefore probes whether an operator can adapt to geometry-induced changes in solution structure when global transport effects play a dominant role. The precise formulation and parameter settings are provided in Supplementary Information A.2.
@@
We adopt the same discrete input–output representation as in the Poisson case (Fig. 2a), consisting of three input channels: geometry mask, boundary-condition locations, and source field; and a single output channel representing the scalar solution pattern. For each transport case, the governing transport parameters are fixed and not treated as additional inputs. The Péclet number ($\mathrm{Pe}$) characterizes the relative strength of advection to diffusion, and separate operators are trained for two representative cases ($\mathrm{Pe}\approx 0.45$ and $\mathrm{Pe}\approx 4.5$). Under this design, increased difficulty arises from the interaction between transport-driven solution structures and geometry-induced discreteness, rather than from changes in the input encoding. It is worth emphasizing that, although the input representation is consistent across different PDEs, the advection–diffusion operator is inherently more complex: its dominant solution patterns are jointly shaped by transport direction, boundary conditions, and geometric constraints, rather than being determined solely by the source distribution.
@@
Figures 3a and b present representative ID validation results for the diffusion-dominated ($\mathrm{Pe}\approx 0.45$) and advection-dominated ($\mathrm{Pe}\approx 4.5$) cases (more results are presented in Supplementary Information E.2). At lower Péclet numbers, the solution patterns remain relatively smooth but already exhibit asymmetric structures aligned with the transport direction. In this case, DiSOL accurately reconstructs the global solution distribution and captures the dominant gradient orientations together with their interaction with geometric boundaries. In contrast, FNO struggles to maintain structural consistency on irregular geometries, producing predictions with noticeable morphological distortions across multiple regions.
@@
As the Péclet number increases, advection becomes the dominant mechanism governing solution behavior. The resulting solution structures exhibit stronger directional transport features and heightened sensitivity to geometric boundaries and local structural variations. Under this advection-dominated setting (Fig. 3b), DiSOL continues to capture the primary transport pathways and their coupling with complex geometries, yielding solution patterns that remain consistent with the reference numerical solutions in both global morphology and key local features. By contrast, FNO displays pronounced structural distortions, with predicted patterns failing to reflect the true transport behavior, particularly near geometric boundaries and internal obstacles.
@@
Taken together, these results demonstrate that, for the advection–diffusion equation governed jointly by geometry and boundary conditions, the limitations of continuous neural operator methods do not stem from increased transport strength alone. Instead, they arise from the amplification of geometry-induced discrete structural variations under transport-dominated conditions. By explicitly modeling local contributions on discrete computational domains and assembling them across multiple scales, DiSOL more naturally captures the coupling between geometry and transport effects, thereby achieving enhanced stability and generalization performance in both ID and OOD settings.

### v2

@@ Abstract
Neural operator learning accelerates PDE solution by approximating operators as mappings between continuous function spaces. Yet in many engineering settings, varying geometry induces discrete structural changes, including topological changes, abrupt changes in boundary conditions or boundary types, and changes in the computational domain, which break the smooth-variation premise. Here we introduce Discrete Solution Operator Learning (DiSOL), a complementary paradigm that learns discrete solution procedures rather than continuous function-space operators. DiSOL factorizes the solver into learnable stages that mirror classical discretizations: local contribution encoding, multiscale assembly, and implicit solution reconstruction on an embedded grid, thereby preserving procedure-level consistency while adapting to geometry-dependent discrete structures. Across geometry-dependent Poisson, advection-diffusion, linear elasticity, as well as spatiotemporal heat conduction problems, DiSOL produces stable and accurate predictions under both in-distribution and strongly out-of-distribution geometries, including discontinuous boundaries and topological changes. These results highlight the need for procedural operator representations in geometry-dominated problems and position discrete solution operator learning as a distinct, complementary direction in scientific machine learning.
@@ 1 Introduction
The numerical solution of partial differential equations (PDEs) lies at the core of scientific computing and engineering analysis [2, 28, 1, 19]. For decades, mature discrete numerical algorithms, such as finite element [28], finite volume [15], and finite difference [30] methods, have provided reliable and accurate solutions across a wide range of applications. A defining characteristic of these methods is that they operate on discrete computational representations: local contributions are evaluated on elements, assembled into global representations, and subsequently solved through well-defined procedures. While geometry variation alters the instantiated discrete system, the underlying solution procedures by which local contributions are evaluated and assembled remain unchanged (Fig. 1a). That is, the same local evaluation and assembly rules are reused wherever they are activated, even though the activated locations change with geometry. In this sense, classical discrete numerical algorithms implement a solution operator: they define a mapping from problem specification (geometry, coefficients, sources, and boundary conditions) to the discrete solution field, executed through an explicit local-to-global procedure.
Recently, data-driven surrogate and operator learning methods have emerged as a promising alternative for accelerating PDE solving across varying inputs and physical parameters [5, 25, 3, 8]. Two complementary directions have been widely explored. One treats PDE learning as field-to-field regression on discrete grids (or strong discretization-based surrogates) using convolutional architectures, often in encoder-decoder or UNet-based forms [22, 21, 36], achieving strong empirical performance when inputs and outputs are naturally represented on fixed-resolution discretizations. The other focuses on continuous neural operators—such as DeepONet [23], the Fourier Neural Operator (FNO) [18], and their variants [9, 12, 14, 16, 35, 17, 10, 24, 32, 31, 7]—that approximate smooth mappings between input and output functions in continuous function spaces (Fig. 1d). Many of these methods have demonstrated impressive performance on benchmark problems with regular domains and smoothly varying geometries [33, 11]. These advances also highlight the role of inductive bias, i.e., structural assumptions encoded in model architectures or learning objectives that shape what solution mechanisms can be learned efficiently from finite data.
@@
This observation highlights a distinction that is often overlooked: although classical discrete numerical solvers can be regarded as operators, learning a continuous operator is not equivalent to learning a discrete numerical solution procedure. Classical solvers do not approximate a single global mapping; instead, they execute a sequence of fixed procedures—local contribution evaluation, global assembly, and solution—which is invariant under geometry variation, even though the assembled global information changes (Figs. 1a,b) [20, 27, 6, 4]. Approximating a smooth functional mapping may therefore be misaligned with the procedural nature of geometry-dependent PDE solution processes.
@@
We validate DiSOL on four classes of geometry-dependent PDEs with increasing complexity: a Poisson equation, an advection-diffusion equation spanning diffusion- and transport-dominated cases, a vector-valued linear elasticity problem, and a spatiotemporal thermal conduction problem. Across these cases, DiSOL delivers stable and accurate predictions under both in-distribution (ID) and out-of-distribution (OOD) computational domains, whereas continuous neural operators exhibit structural degradation when geometry-induced discrete variation dominates the problem setting. Moreover, our results indicate that these failures arise primarily from a mismatch between the assumptions of continuous operators and geometry-induced discrete solution structures, rather than from optimization difficulty or insufficient model capacity.
@@
Fig. 1: Discrete Solution Operator Learning (DiSOL) in contrast to continuous neural operator paradigms. a, Classical discrete numerical methods for partial differential equations (PDEs), exemplified by the finite element method. A PDE operator can be realized through a sequence of fixed discrete procedures, including local operator evaluation (e.g., element-level stiffness and force computation), global assembly, and solution via assembled global systems. Under geometry variation, the assembled results change, whereas the underlying assembly procedures remain unchanged. b, An abstracted view of discrete problem solving, highlighting the canonical workflow shared by many numerical methods: local contributions evaluation, global information assembly, and solution through globally aggregated representations. c, Discrete Solution Operator Learning (DiSOL). Instead of explicitly constructing numerical matrices, DiSOL represents the discrete solution process through learnable neural operators that conceptually correspond to the local operator, the multiscale assembly operator, and the implicit problem-solving operator. These components collectively form a discrete operator that preserves procedure-level invariance under geometry variation, in the sense that the underlying solution procedure remains unchanged, while adapting its outputs to geometry-dependent discrete structures. d, Representative continuous neural operator paradigms, illustrated by DeepONet [23] and FNO [18]. These approaches model PDE operators as global mappings in continuous function spaces, using global basis combinations or spectral transformations. Such representations are primarily designed for smooth function-space mappings, and may become misaligned when geometry changes and discontinuities dominate the problem setting. In contrast, DiSOL does not aim to approximate continuous operators directly, but learns discrete solution procedures, leading to a fundamentally different inductive bias in geometry-dependent problem settings.
@@ 2 Results
where $\Omega_{h}$ denotes the discrete computational domain (the geometry mask and the fixed grid-neighborhood connectivity), $\Gamma_{h}$ encodes boundary-type on that domain, $f_{h}$ denotes source fields, $\eta_{h}$ collects problem-dependent physical parameters (e.g., diffusion coefficient, advection velocity, elastic moduli) that are fixed or prescribed for a given task, and $u_{h}$ denotes the normalized solution pattern field on the grid. In the main text we focus on learning the normalized solution pattern $u_{h}:=U_{h}/u_{\lim}$ (see Methods, Section 4.1 and Supplementary Information D); all models predict $\hat{u}_{h}$, and all reported metrics are computed on $u_{h}$ restricted to $\Omega_{h}$. All experiments below follow Eq. 1.
@@ 2.1 Geometry-dependent Poisson problem
Fig. 2: Discrete solution operator learning (DiSOL) for a geometry-dependent 2D Poisson problem. a, Problem formulation and data representation. The input to the model consists of three discrete fields defined on a fixed Cartesian grid: a geometry mask, a boundary-condition selection map, and a source term. The model outputs the corresponding discrete solution pattern. b, Training and validation loss histories for DiSOL, DeepONet and the FNO under comparable model capacity and training settings. DiSOL converges faster and achieves substantially lower validation loss (All models are trained under the same optimizer/schedule and have comparable parameter counts, $\approx 0.13$M each; details in Supplementary Information C.4 and C.5). c, Representative ID validation cases. While the reference solutions are globally smooth and continuous operator baselines capture the coarse trend, DeepONet and FNO show mild yet systematic deviations in the normalized pattern, particularly near irregular geometry and boundary-adjacent regions where geometry-conditioned boundary activation influences local responses. DiSOL provides closer agreement with the ground truth and achieves lower errors under the same ID protocol. d, OOD generalization tests involving unseen geometries with sharp corners and internal holes, discontinuous and localized boundary conditions, and high-frequency source terms. DiSOL remains stable and consistent with the reference solutions, whereas DeepONet and FNO show substantial degradation in solution-pattern fidelity. See Supplementary Information A.1, B and E.1 for full setup and extended statistics.
@@
Figure 2c presents two representative ID test cases (more results are presented in Supplementary Information E.1). Since these samples follow the training distribution and yield globally smooth Poisson solutions, both continuous-operator baselines are able to capture the coarse solution trend. Nevertheless, compared with DiSOL, DeepONet (and FNO) exhibit consistent residual discrepancies in the normalized pattern, which are most visible near irregular geometric features and boundary-adjacent regions, where the solution is jointly shaped by the discrete activation of boundary segments and the local domain geometry. These deviations manifest as mild distortions in the relative distribution of values and weakened boundary-conditioned contrast, rather than random noise. DiSOL more faithfully reproduces both the global morphology and the boundary-modulated local variations, leading to noticeably lower errors under the same ID setting. A capacity-matched U-Net control study is provided in Supplementary Information H to disentangle backbone effects from the discrete-operator formulation.
@@
2.2 Geometry-dependent advection-diffusion problem
@@
Fig. 3: Discrete solution operator learning (DiSOL) results for the advection-diffusion equation. a,b, Representative ID validation cases under two transport settings: a diffusion-dominated case ($\text{Pe}\approx 0.45$) and an advection-dominated case ($\text{Pe}\approx 4.5$). Ground-truth solutions are compared with predictions from DiSOL and FNO. As transport effects become stronger and solution structures align more closely with flow direction and complex geometric boundaries, DiSOL consistently preserves the global transport patterns and boundary-induced features, whereas FNO exhibits increasingly pronounced structural deviations, particularly near irregular geometric features. c,d, Training and validation loss histories for DiSOL, FNO, and DeepONet under the same two Péclet number settings ($\text{Pe}\approx 0.45$ and $\text{Pe}\approx 4.5$). In both cases, DiSOL converges rapidly and achieves substantially lower training and validation errors, while the continuous neural operator baselines saturate at significantly higher error levels. e, OOD generalization results for the advection-diffusion equation at $\text{Pe}\approx 4.5$, involving unseen geometries with sharp corners and topological changes, together with localized and discontinuous boundary conditions. DiSOL preserves the overall transport structure and solution morphology. f, Statistical comparison of relative L1 errors across different models under ID and OOD settings for both $\text{Pe}\approx 0.45$ and $\text{Pe}\approx 4.5$. Box plots (logarithmic scale) show that DiSOL consistently outperforms continuous neural operator baselines in both cases, and that performance degradation is primarily driven by geometric distribution shifts rather than by increased transport dominance. See Supplementary Information A.2, B and E.2 for full setup and extended statistics.
@@
We next consider a steady advection-diffusion problem to examine geometry generalization under transport-dominated dynamics. Although the governing equation and physical parameters are fixed within each case, variations in domain geometry induce sharp changes in effective transport pathways and boundary-layer structures. This case therefore probes whether an operator can adapt to geometry-induced changes in solution structure when global transport effects play a dominant role. The precise formulation and parameter settings are provided in Supplementary Information A.2.
@@
We adopt the same discrete input-output representation as in the Poisson case (Fig. 2a), consisting of three input channels: geometry mask, boundary-condition locations, and source field; and a single output channel representing the scalar solution pattern. In this problem, the Péclet number ($\mathrm{Pe}$) characterizes the relative strength of advection to diffusion [13], and separate operators are trained for two representative cases ($\mathrm{Pe}\approx 0.45$ and $\mathrm{Pe}\approx 4.5$). Under this design, increased difficulty arises from the interaction between transport-driven solution structures and geometry-induced discreteness, rather than from changes in the input encoding. It is worth emphasizing that, although the input representation is consistent across different PDEs, the advection-diffusion operator is inherently more complex: its dominant solution patterns are jointly shaped by transport direction, boundary conditions, and geometric constraints, rather than being determined solely by the source distribution.
@@
Figures 3a and b present representative ID validation results for the diffusion-dominated ($\mathrm{Pe}\approx 0.45$) and advection-dominated ($\mathrm{Pe}\approx 4.5$) cases (more results are presented in Supplementary Information E.2). At lower Péclet numbers, the solution patterns remain relatively smooth but already exhibit asymmetric structures aligned with the transport direction. As observed, DiSOL accurately predicts the global solution patterns with their randomly selected boundaries. In contrast, FNO struggles to maintain structural consistency on irregular geometries, producing predictions with noticeable morphological distortions across multiple regions.
@@
As the Péclet number increases, advection becomes the dominant mechanism governing solution behavior. The resulting solution structures exhibit stronger directional transport features and heightened sensitivity to geometric boundaries and local structural variations. Under this advection-dominated setting (Fig. 3b), DiSOL continues to capture the primary transport pathways and their coupling with complex geometries, yielding solution patterns that remain consistent with the reference numerical solutions in both global morphology and key local features. By contrast, FNO displays pronounced structural distortions, with predicted patterns failing to reflect the true transport behavior, particularly near geometric boundaries.
@@
Taken together, these results demonstrate that, for the advection-diffusion equation governed jointly by geometry and boundary conditions, the limitations of continuous neural operator methods do not stem from increased transport strength alone. Instead, they arise because transport-dominated dynamics are highly sensitive to geometry-induced discrete changes, which can markedly alter the solution. By explicitly modeling local contributions on discrete computational domains and assembling them across multiple scales, DiSOL more naturally captures the coupling between geometry and transport effects, thereby achieving enhanced stability and generalization performance in both ID and OOD settings.